Write the function to show a collectable doll that has doubled in value in 10 years. The initial value of the doll was $355.00.
I think it is y=355(2)^x/10 ???????????????

Answer by edjones(8007)

355*10x=710
10x=710/355
10x=2
x=2/10
x=.2
f(.2)=355*10x
Ed